Brown Industries has a debt-equity ratio of 1.6. Its WACC is 10 percent, and its cost of debt is 7 percent. There is no corporate tax.

What would the cost of equity be if the debt-equity ratio were 2? (Do not round intermediate calculations and enter your answer as a percent rounded to the nearest whole number, e.g., 32.)

What is the company’s cost of equity capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

What would the cost of equity be if the debt-equity ratio were .4? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

What would the cost of equity be if the debt-equity ratio were zero? (Do not round intermediate calculations and enter your answer as a percent rounded to the nearest whole number, e.g., 32.)
Answer #1


WACC=rd*(1-t)*wd+rs *ws 10=7*(1-0)*(1.6/2.6)+rs* 1/2.6 10-7*(1-0)*(1.6/2.6)=rs*1/2.6 rs=(10-7*(1-0) *(1.6/2.6))/(1/2.6) rss 1WACC=rd*(1-t)*wd+rs *ws 10=7*(1-0) * (2/3)+rs*1/3 10-7*(1-0)*(2/3)=rs* 1/3 rs=(10-7*(1-0)*(2/3)/(1/3) rs 16.00%WACC=rd*(1-t)*wd+rs *ws 10=7*(1-0)*(0.4/1.4)+rs*1/1.4 10-7*(1-0)*(0.4/1.4)=rs *1/1.4 rs=10-7*(1-0)*(0.4/1.4))/(1/1.4) rs 11.WACC=rd*(1-t)*wd+rs *ws 10=7*(1-0)*(0/1)+ rs 1/1 10-7*(1-0)*(0/1)=rs* 1/1 rs=10-7*(1-0)*(0/1))/(1/1) 10.00%